# Diffusion Model as a Noise-Aware Latent Reward Model for Step-Level Preference Optimization

### v2

@@ 5.2 Main Results
Quantitative Comparison. As indicated in Tab. 2, Tab. 3, and Tab. 5, Diffusion-DPO excels in enhancing the text-image alignment, while SPO focuses more on aesthetics. LPO outperforms both methods across three dimensions, achieving higher Aesthetic Scores and superior performance on T2I-CompBench++ and GenEval metrics, leading to improved general preference scores. The user study results indicate similar findings, as discussed in Appendix B. Notably, the LPO-optimized SD1.5 even exhibits performance comparable to the original SDXL model across various metrics. We further validate the effectiveness of heterogeneous optimization in Tab. 6. SD1.5 and SD2.1 (Rombach et al., 2022) share the same VAE encoder, but SD1.5 has a smaller text encoder. Remarkably, fine-tuning SD2.1 using LRM-1.5 still yields significant improvements across various aspects, demonstrating that a smaller and inferior diffusion model can effectively fine-tune a larger and more advanced model as long as they share the same VAE encoder. In contrast, applying LRM-1.5 for the LPO of SDXL is ineffective due to the distribution mismatch in their latent spaces, which arises from differences in their VAE encoders.
Qualitative Comparison. The qualitative comparisons of various methods are illustrated in Fig. 2 and Fig. 11-Fig. 15. The images generated by Diffusion-DPO exhibit deficiencies in color and detail, whereas those produced by SPO demonstrate lower semantic relevance. Additionally, SPO’s excessive focus on aesthetics may lead to an overabundance of details in some images, making them appear cluttered. In contrast, the images produced by LPO achieve a strong balance between text-image alignment and aesthetic quality, delivering a higher overall image quality.
Training Efficiency Comparison. LPO achieves significantly faster training speed. As shown in Tab. 5, considering the time required for both reward modeling and preference optimization, LPO requires only 23 A100 hours for SD1.5—just 1/10 of the training time needed for Diffusion-DPO and 1/3.5 of that for SPO. For SDXL, LPO’s training time is reduced to 1/28 and 1/2.5 of that for Diffusion-DPO and SPO, respectively. This efficiency is primarily due to LPO performing reward modeling and preference optimization directly in the latent space, avoiding the additional computational overhead of converting to pixel space.
@@ 5.3 Ablation Studies
If not specified, ablation experiments are conducted on SD1.5. Due to space limitations, we only use PickScore to reflect general preference in Tab. 7 and Tab. 8.
MPCF. As shown in Tab. 7, MPCF plays a critical role in LRM training. As discussed in Sec. 4.2, the inconsistent preference issue makes training on the full dataset (wo MPCF) ineffective, since it hinders the LRM from adequately focusing on aesthetics or text-image alignment, resulting in inferior LPO performance. On the other hand, different filtering strategies can profoundly impact the preference patterns of both the LRM and LPO-optimized models. The first filtering strategy strictly requires that winning images score higher than losing images across all aspects. However, since the diffusion model lacks explicit text-image alignment pre-training like CLIP, it is prone to overfitting to the visual features of the images, as indicated by a higher Aes-Corr. This overfitting results in reduced attention to alignment, as reflected by lower CLIP-Corr and GenEval scores. The second and third strategies relax the aesthetic constraints to varying degrees. However, excessively lenient constraints (the 3rd strategy) may cause LRM to focus solely on text-image alignment while neglecting image quality, resulting in a negative Aes-Corr. In contrast, the second strategy balances these two aspects better, leading to the highest general preference scores.
Table 7: Ablation results on MPCF of LRM’s training data. The second strategy balances aesthetics and alignment better.
@@
Table 8: Ablation results on the VFE module of LRM. Introducing VFE leads to better alignment and general preferences.
@@
Table 9: Ablation results on optimization timestep ranges in LPO.
Range of $t$ Aesthetic GenEval P-S I-R HPSv2 HPSv2.1
[0, 200] 5.434 40.11 20.46 -0.0987 26.25 23.61
[250, 450] 5.527 43.00 20.76 0.1430 26.90 25.37
[500, 700] 5.742 44.44 20.95 0.1591 26.71 25.16
[750, 950] 5.853 48.28 21.54 0.6337 27.47 27.64
[0, 450] 5.573 42.71 20.63 0.0204 26.69 24.88
[0, 700] 5.765 44.93 21.02 0.3087 27.10 26.25
[0, 950] 5.945 48.39 21.69 0.6588 27.64 27.86
Table 10: Ablation results on different threshold strategies.
Threshold Aesthetic GenEval P-S I-R HPSv2 HPSv2.1
0.3 5.853 46.75 21.22 0.5112 27.30 27.12
0.4 5.832 48.32 21.32 0.4789 27.08 26.37
0.5 5.900 48.39 21.57 0.6088 27.54 27.42
0.6 5.877 47.97 21.35 0.5510 27.25 26.73
[0.3, 0.45] 5.916 49.43 21.58 0.6405 27.55 27.33
[0.35, 0.5] 5.945 48.39 21.69 0.6588 27.64 27.86
[0.4, 0.55] 5.882 48.77 21.48 0.4791 27.30 27.13
Structure of LRM. As illustrated in Tab. 8, the introduction of VFE ($gs>1$) leads to lower Aes-Corr values but higher CLIP-Corr values, indicating an enhanced emphasis on text-image alignment. This results in improvements in both the GenEval score and PickScore, with only a minor decline in the Aesthetic Score. As $gs$ increases, the LRM’s correlation with alignment steadily improves, while its correlation with aesthetics decreases. When $gs$ is set to 7.5, the model achieves the best overall performance.
Optimization Timesteps. Tab. 9 ablates different optimization timestep ranges, indicating that larger timesteps lead to better performance. The results achieved within the range of $[750,950]$ are nearly comparable to those achieved through optimization across the entire denoising process, i.e. $[0,950]$. We suggest this is because diffusion models focus on low-frequency information, such as image layout and style, during larger timesteps, while emphasizing high-frequency texture details during smaller timesteps. The low-frequency components formed in higher timesteps play a decisive role in determining the overall quality of the generated images. This observation also demonstrates the effectiveness of LRM, even in very large timesteps. The qualitative comparison of different ranges is shown in Fig. 10.
Dynamic Sampling Threshold. The standard deviation $\sigma_{t}$ of samples at smaller timesteps is relatively small according to the DDPM scheduling (Ho et al., 2020), making the constant threshold insufficient to accommodate all timesteps. As indicated in Tab. 10, the dynamic threshold strategy generally outperforms the constant threshold across different intervals, effectively alleviating this problem. We further explore other dynamic strategies in Appendix B.
@@ 6 Discussion and Conclusion
Conclusion. In this paper, we propose LRM to utilize diffusion models for step-level reward modeling, based on the insights that diffusion models possess text-image alignment abilities and can perceive noisy latent images across different timesteps. To facilitate the training of LRM, the MPCF strategy is introduced to address the inconsistent preference issue in LRM’s training data. We further propose LPO, a method that employs LRM for step-level preference optimization, operating entirely within the latent space. LPO not only significantly reduces training time but also delivers remarkable performance improvements across various evaluation dimensions, highlighting the effectiveness of employing the diffusion model itself to guide its preference optimization. We hope our findings can open new avenues for research in preference optimization for diffusion models and contribute to advancing the field of visual generation.
Limitations and Future Work. (1) The experiments in this work are conducted on UNet-based models and the DDPM scheduling method. Further research is needed to adapt these findings to larger DiT-based models (Esser et al., 2024) and flow matching methods (Lipman et al., 2023). (2) The Pick-a-Pic dataset mainly contains images generated by SD1.5 and SDXL, which generally exhibit low image quality. Introducing higher-quality images is expected to enhance the generalization of the LRM. (3) As a step-level reward model, the LRM can be easily applied to reward fine-tuning methods (Prabhudesai et al., 2023; Clark et al., 2024), avoiding lengthy inference chain backpropagation and significantly accelerating the training speed. (4) The LRM can also extend the best-of-N approach to a step-level version, enabling exploration and selection at each step of image generation, thereby achieving inference-time optimization similar to GPT-o1 (gpt, 2024).
Impact Statement
This work introduces a preference optimization method used for text-to-image diffusion models. The method may have the following impacts:
@@ Appendix A Experimantal Details
LRM Training. We train LRM-1.5 for 4,000 steps with 500 warmup steps and LRM-XL for 8,000 steps with 1000 warmup steps, both using the learning rate 1e-5. The batch size is 128 for LRM-1.5 and 32 for LRM-XL, respectively. Following CLIP (Radford et al., 2021), the initial value of $\tau$ in Eqn (10) is set to $e^{2.6592}$. The training resolution is 512 for SD1.5 and 1024 for SDXL. For SDXL, which includes two text encoders, we utilize only the OpenCLIP ViT-bigG (Cherti et al., 2023) as the text encoder of LRM-XL.
LPO Training. Both SD1.5 and SDXL are fine-tuned for 5 epochs. Following SPO (Liang et al., 2024), we employ LoRA (Hu et al., 2022) to fine-tune models. The LoRA rank is 4 for SD1.5 and 64 for SDXL. All experiments are conducted on 4 A100 GPU. Other hyperparameters are provided in Tab. 11. The experimental setting of SD2.1 is the same as SD1.5.
@@
Baseline Methods. Specifically, SPO and LPO are trained on Pick-a-Pic v1 (Kirstain et al., 2023) while Diffusion-DPO and MaPO are trained on Pick-a-Pic v2 (Kirstain et al., 2023), an extended version of Pick-a-Pic v1. Pick-a-Pic v1 consists of 583,737 preference pairs, whereas Pick-a-Pic v2 contains over 950,000 pairs. In the practice of SPO, DDPO (Black et al., 2024) and D3PO (Yang et al., 2024) are reproduced on 4k prompts randomly sampled from Pick-a-Pic v1, which are the same as the training data used for SPO and LPO. Therefore, the comparison between these methods can be considered fair.
Appendix B Additional Experiments
@@
Figure 6: User study results on 200 prompts from HPSv2 benchmark (Wu et al., 2023) and DPG-Bench (Hu et al., 2024).
@@
User Study. We conduct human evaluation experiments for LPO, SPO, and Diffusion-DPO. The evaluation set consists of 200 prompts, with 100 randomly sampled from HPSv2 benchmark (Wu et al., 2023) and 100 randomly sampled from DPG-Bench (Hu et al., 2024). For each prompt, the model generates four images. For each image, five expert evaluators score it across three dimensions: general preference, visual appeal, and prompt alignment. Subsequently, votes are cast to determine the winning relationships among the different models. The results are illustrated in Fig. 6. Similar to the results of automatic metrics in the main paper, LPO outperforms both SPO and Diffusion-DPO across three dimensions. Compared to T2I-CompBench and GenEval results, the advantage of LPO on text-image alignment in the human evaluation results is not particularly pronounced. We believe this is because the prompts used for human evaluation are not specifically designed to assess text-image alignment, and different models exhibit relatively good alignment.
@@
Specific Reward Model (VLM-Based) Aesthetic (Schuhmann et al., 2022) - - - 54.03
CLIP Score (CLIP-H) (Radford et al., 2021) - - - 61.84
VQAScore (Lin et al., 2025) - - - 59.16
General Reward Model (VLM-Based) ImageReward (Xu et al., 2023) 0.108 0.425 0.417 62.66
HPSv2 (Wu et al., 2023) 0.007 0.602 0.406 64.76
HPSv2.1 (Wu et al., 2023) 0.191 0.432 0.332 65.58
PickScore (Kirstain et al., 2023) 0.066 0.490 0.402 71.93
@@
LRM as Pixel-Wise Reward Model. The LRM can also serve as a pixel-wise reward model when combined with the corresponding VAE encoder. Tab. 12 shows the correlations of different reward models with aesthetics and text-image alignment. Aes-Corr is employed to assess the correlation with aesthetics, while the CLIP-Corr and VQA-Corr are utilized to measure alignment correlation. The calculation method for VQA-Corr is similar to that of CLIP-Corr, as described in Sec. 5.1, with the CLIP Score replaced by VQAScore. It is observed that the LRMs, especially LRM-XL, exhibit aesthetic and alignment correlations comparable to those of VLM-based models, as indicated by similar Aes-Corr and VQA-Corr values. The CLIP-Corr of LRMs is slightly lower than the average value of VLM-based reward models, likely because VLM-based models are fine-tuned versions of CLIP (Radford et al., 2021) or BLIP (Li et al., 2022), leading to greater similarity to CLIP. The accuracy on the validation and test sets of Pick-a-Pic is also present in Tab. 12. LRMs demonstrate competitive performance, with larger models (LRM-XL) exhibiting higher accuracy. Moreover, we find that adding slight noise to the latent images ($t=200$) can achieve better accuracy. This can be attributed to the fact that slight noise can better activate the feature extraction capabilities and improve the generalization of diffusion models.

### v3

@@ 5.2 Main Results
Quantitative Comparison. As indicated in Tab. 2 and Tab. 3, LPO outperforms other methods across various dimensions by a large margin on SD1.5, especially in general preference and text-image alignment. Based on SDXL, LPO even slightly surpasses InterComp, although the latter utilizes a superior dataset comprising images from SD3[1] and FLUX [43]. The user study results in Appendix B further confirm the superiority of LPO. We also validate the effectiveness of heterogeneous optimization in Tab. 6. Remarkably, fine-tuning SD2.1 using LRM-1.5 yields significant improvements across various aspects, demonstrating that an inferior diffusion model can effectively fine-tune an advanced model as long as they share the same VAE encoder. In contrast, applying LRM-1.5 for the LPO of SDXL is ineffective due to a distribution mismatch between the latent spaces of their VAE encoders.
Qualitative Comparison. The qualitative comparisons of various methods are illustrated in Fig. 1 and Fig. 11-Fig. 15. Images generated by Diffusion-DPO exhibit deficiencies in color and detail, whereas those from SPO demonstrate reduced semantic relevance and excessive details in some images, resulting in cluttered visuals. In contrast, the images produced by LPO achieve a strong balance between text-image alignment and aesthetic quality, delivering a higher overall image quality.
Training Efficiency Comparison. LPO enables significantly faster training. As shown in Tab. 6, by performing reward modeling and preference optimization directly in the noisy latent space, LRM bypasses both $\hat{x}_{0,t}$ prediction and VAE decoding during sampling. When $K=4$, its sampling time is only 1/6 that of SPM in SPO, yielding substantial time savings—and the same efficiency gain applies to LRM training. Consequently, as illustrated in Tab. 6, LPO requires only 23 A100 hours for SD1.5—just 1/10 the training time of Diffusion-DPO and 1/3.5 that of SPO. On SDXL, LPO’s training time is reduced to 1/28 and 1/2.5 of that for Diffusion-DPO and SPO, respectively.
Further Explorations. In Appendix B, we present additional exploration experiments on LRM and LPO. First, we validate the effectiveness of LRM on a step-wise variant of GRPO [15]. We then assess the applicability of LPO on DiT-Based SD3 [1] and evaluate its generalization on other datasets. Additionally, we report the performance of LRM when used as a pixel-wise reward model.
@@ 5.3 Ablation Studies
MPCF. As shown in Tab. 8, MPCF plays a vital role in LRM training. The first filtering strategy enforces that winning images must outperform losing images across all aspects. However, since the diffusion model lacks explicit text-image alignment pre-training like CLIP, it is prone to overfitting to visual features, as indicated by a higher Aes-Corr. This overfitting results in reduced attention to alignment, as reflected by lower CLIP-Corr and GenEval scores. The second and third strategies relax the aesthetic constraints to varying degrees. While the third, most lenient strategy can cause LRM to focus solely on alignment and neglect image quality, as evidenced by a negative Aes-Corr value, the second strategy achieves a better balance, yielding the highest general preference scores. Notably, even without MPCF and using the same training data, LPO still outperforms SPO in both general and text-image alignment preferences, as detailed in Appendix C.
Table 7: Ablation results on LRM’s training data.
@@
Table 8: Ablation results on the VFE module.
@@
Table 9: Ablation results on timestep ranges.
Range of $t$ P-S I-R HPSv2 HPSv2.1 Aesthetic GenEval
[0, 200] 20.46 -0.0987 26.25 23.61 5.434 40.11
[250, 450] 20.76 0.1430 26.90 25.37 5.527 43.00
[500, 700] 20.95 0.1591 26.71 25.16 5.742 44.44
[750, 950] 21.54 0.6337 27.47 27.64 5.853 48.28
[0, 450] 20.63 0.0204 26.69 24.88 5.573 42.71
[0, 700] 21.02 0.3087 27.10 26.25 5.765 44.93
[0, 950] 21.69 0.6588 27.64 27.86 5.945 48.39
Table 10: Ablation results on threshold strategies.
Threshold P-S I-R HPSv2 HPSv2.1 Aesthetic GenEval
0.3 21.22 0.5112 27.30 27.12 5.853 46.75
0.4 21.32 0.4789 27.08 26.37 5.832 48.32
0.5 21.57 0.6088 27.54 27.42 5.900 48.39
0.6 21.35 0.5510 27.25 26.73 5.877 47.97
[0.3, 0.45] 21.58 0.6405 27.55 27.33 5.916 49.43
[0.35, 0.5] 21.69 0.6588 27.64 27.86 5.945 48.39
[0.4, 0.55] 21.48 0.4791 27.30 27.13 5.882 48.77
Structure of LRM. As illustrated in Tab. 8, the introduction of VFE ($gs>1$) leads to lower Aes-Corr values but higher CLIP-Corr values, indicating an enhanced emphasis on text-image alignment. This results in improvements in both the GenEval and PickScore, with only a minor decline in Aesthetic Score. As $gs$ increases, the LRM’s correlation with alignment steadily improves, while its correlation with aesthetics decreases. When $gs$ is set to 7.5, the model achieves the best overall performance.
Optimization Timesteps. Tab. 10 ablates different optimization timestep ranges, indicating that broader ranges lead to better performance. Notably, LRM can effectively predict preferences even at very large timesteps, e.g., $[750,950]$, achieving results comparable to those obtained across the entire denoising process, i.e. $[0,950]$. This highlights the critical role of large timesteps in step-level preference optimization. In contrast, SPO performs poorly in large timesteps, as indicated in Tab. 23 in Appendix C. This comparison underscores the superiority of using diffusion models for step-level reward modeling at large timesteps, since LRM can directly process noisy latent images without suffering from distribution shift. A qualitative comparison of different ranges is provided in Fig. 10.
Dynamic Sampling Threshold. The standard deviation $\sigma_{t}$ of samples at smaller timesteps is relatively small according to the DDPM scheduling [44], making the constant threshold insufficient to accommodate all timesteps. As indicated in Tab. 10, the dynamic threshold strategy generally outperforms the constant threshold across different intervals, effectively alleviating this problem. We further explore other dynamic strategies in Tab. 26 in Appendix C.
@@ 6 Discussion and Conclusion
Conclusion. In this paper, we propose LRM, a method that utilizes diffusion models for step-level reward modeling directly in the noisy latent space, based on the insights that diffusion models inherently possess text-image alignment abilities and can effectively perceive noisy latent images across different timesteps. To facilitate LRM training, we introduce the MPCF strategy to mitigate the inconsistent preference issue in training data. Building on LRM, we further propose LPO, a step-level preference optimization that operates entirely within the noisy latent space. LPO not only achieves substantial training speedups but also delivers remarkable performance improvements across various evaluation metrics, highlighting the effectiveness of employing the diffusion model itself to guide its preference optimization. We hope our findings can open new avenues for research in preference optimization for diffusion models and contribute to advancing the field of visual generation.
Limitations and Future Work. (1) The Pick-a-Pic dataset mainly contains images generated by SD1.5 and SDXL, which generally exhibit low image quality. Introducing higher-quality images is expected to enhance the generalization of the LRM. (2) As a step-level reward model, the LRM can be easily applied to reward fine-tuning methods [45, 24, 26], avoiding lengthy inference chain backpropagation and significantly accelerating the training speed. (3) The LRM can also extend the best-of-N approach to a step-level version, enabling exploration and selection at each step of image generation, thereby achieving inference-time optimization similar to GPT-o1 [46].
@@ Appendix A Experimantal Details
LRM Training. We train LRM-1.5 for 4,000 steps with 500 warmup steps and LRM-XL for 8,000 steps with 1000 warmup steps, both using the learning rate 1e-5. The batch size is 128 for LRM-1.5 and 32 for LRM-XL, respectively. Following CLIP [9], the initial value of $\tau$ in Eqn. (9) is set to $e^{2.6592}$. The training resolution is 512 for SD1.5 and 1024 for SDXL. For SDXL, which includes two text encoders, we utilize only the OpenCLIP ViT-bigG [47] as the text encoder of LRM-XL.
LPO Training. Both SD1.5 and SDXL are fine-tuned for 5 epochs. Following SPO [8], we employ LoRA [48] to fine-tune models. The LoRA rank is 4 for SD1.5 and 64 for SDXL. All experiments are conducted using 4 A100. Other hyperparameters are provided in Tab. 11. The experimental setting of SD2.1 is the same as SD1.5.
@@
Baseline Methods. Specifically, SPO [8] and LPO are trained on Pick-a-Pic v1 [16] while Diffusion-DPO [4], SePPO [30], and MaPO [29] are trained on Pick-a-Pic v2 [16], an extended version of Pick-a-Pic v1. Pick-a-Pic v1 consists of 583,737 preference pairs, whereas Pick-a-Pic v2 contains over 950,000 pairs. In the practice of SPO, DDPO [6] and D3PO [7] are reproduced on 4k prompts randomly sampled from Pick-a-Pic v1, which are the same as the training data used for SPO and LPO. Therefore, the comparison between these methods can be considered fair. Furthermore, for reference purposes, we also make a comparison with InterComp [31]. Instead of Pick-a-Pic, InterComp utilizes an internally constructed preference dataset of higher quality, including images generated by more advanced models such as SD3 [1] and FLUX [43]. In contrast, the data in Pick-a-Pic only consists of images from models like SDXL [11] and SD2.1 [14]. Despite the relatively low-quality training data, LPO still performs slightly better than InterComp, demonstrating the effectiveness of our method.
Evaluation Practice. For all baseline methods, we use the official models for evaluation without retraining. For the evaluation of text-image alignment, we utilize official codebases of GenEval [40] and T2I-CompBench++ [41]. Following SPO, we employ several public preference models to assess general and aesthetic preferences, including PickScore [16], ImageReward [18], HPSv2 [17], HPSv2.1 [17], and Aesthetic Predictor [38]. Since SPO does not provide the evaluation code, we implement our pipeline, adhering to SPO’s setting: using the DDIM [35] scheduler with 20 inference steps and the same random seed to ensure reproducibility. We apply the same code for all models to ensure a fair comparison. Although there are some discrepancies with SPO’s reported results, we believe the relative comparisons between different models are reliable and meaningful.
Appendix B Additional Exploration Experiments
@@
Figure 6: User study results on 200 prompts from HPSv2 benchmark [17] and DPG-Bench [49].
@@
User Study. We conduct human evaluation experiments for LPO, SPO, and Diffusion-DPO. The evaluation set consists of 200 prompts, with 100 randomly sampled from HPSv2 benchmark [17] and 100 randomly sampled from DPG-Bench [49]. For each prompt, the model generates four images. For each image, five expert evaluators score it across three dimensions: general preference, visual appeal, and prompt alignment. Subsequently, votes are cast to determine the winning relationships among the different models. The results are illustrated in Fig. 6. Similar to the results of automatic metrics in the main paper, LPO outperforms both SPO and Diffusion-DPO across three dimensions. Compared to T2I-CompBench and GenEval results, the advantage of LPO on text-image alignment in the human evaluation results is not particularly pronounced. We believe this is because the prompts used for human evaluation are not specifically designed to assess text-image alignment, and different models exhibit relatively good alignment.
Exploration of GRPO Based on LRM. Actually, LRM is not strictly bound to LPO, and it can also be applied to other preference optimization methods. Inspired by DeepSeekMath [15], we implement a step-wise GRPO algorithm for preference optimization in the image domain and use LRM as the reward model. Specifically, at each denoising step from $x_{t+1}$ to $x_{t}$, we sample a group of $x_{t}^{i},i=1,...,K$ according to Eqn. (2). The LRM is employed to provide rewards $R_{t}=\{r_{t}^{i},i=1,...,K\}$ for these noisy latent images. We then normalize these rewards to calculate corresponding advantages by $A_{t}^{i}=\frac{r_{t}^{i}-mean(R_{t})}{std(R_{t})}$. Finally, we optimize diffusion models by minimizing the following objective:
$L_{GRPO}=-\mathbb{E}[{\{x_{t}^{i}\}_{i=1}^{K}\sim p_{\theta}(x_{t}|x_{t+1},c)}]$ (13)
$\left\{\min\left[\frac{p_{\theta}(x_{t}^{i}|x_{t+1},c)}{p_{\theta_{old}}(x_{t}^{i}|x_{t+1},c)},\text{clip}\left(\frac{p_{\theta}(x_{t}^{i}|x_{t+1},c)}{p_{\theta_{old}}(x_{t}^{i}|x_{t+1},c)},1-\epsilon,1+\epsilon\right)\right]A_{t}^{i}-\beta\mathbb{D}_{KL}[p_{\theta}||p_{ref}]\right\}$
We conduct preliminary experiments of step-wise GRPO on the SD1.5 model. We set $K$ to 4, $\beta$ to 0.1 and $\epsilon$ to 0.1. As shown in Tab. 14, based on our LRM-1.5, step-wise GRPO consistently improves the model’s performance across various dimensions. However, in the current setting, step-wise GRPO performs slightly worse than LPO and requires 8$\times$ the training time. We will further explore the application of LRM-based GRPO for image-domain preference optimization in future work.
Table 12: The performance of different preference optimization methods based on LRM. The general and aesthetic scores are calculated on the Pick-a-Pic validation unique set. We use 20 inference steps.
Model Method PickScore ImageReward HPSv2 HPSv2.1 Aesthetic GenEval (20 Step) Training Time $\downarrow$ (A100 hour)
SD1.5 [14] Original 20.56 0.0076 26.46 24.05 5.468 42.56 -
LPO 21.69 0.6588 27.64 27.86 5.945 48.39 8
Step-Wise GRPO 21.52 0.5861 27.47 28.35 6.087 45.88 64
Table 13: General and aesthetic preference scores on Pick-a-Pic validation unique set, along with GenEval [40] scores. We use 20 inference steps.
Model Method Dataset PickScore ImageReward HPSv2 HPSv2.1 Aesthetic
SD3-Medium [1] Original - 21.74 0.9527 28.27 28.74 5.741
LPO Pick-a-Pic v1 [16] 22.29 1.1306 28.64 30.44 5.993
Table 14: Qualitative results on GenEval [40] with 20 inference steps based on SD3-Medium.
SD3-Medium [1] Original 98.13 81.57 57.81 82.98 24.50 58.75 67.29
LPO 100.00 86.11 60.63 84.04 26.25 54.75 68.63
Exploration on DiT-Based Models with Flow Matching Methods (SD3). In this paper, we mainly conduct experiments on U-Net-based models with the DDPM [44] scheduling method. Here we explore the effectiveness of LRM and LPO on DiT-based models with the Flow Matching method [50]. Firstly, based on SD3-medium [1], we train the LRM on Pick-a-Pic v1 [16]. We extract visual features from the 18th layer of MMDiT [1] and utilize text features from all text encoders. The encoders of the SD-medium are frozen during training. Secondly, to provide the exploration space for LPO, we transform Flow Matching from ODE (Ordinary Differential Equation) to SDE (Stochastic Differential Equation). Specifically, we replace predicted noise $\epsilon_{\theta,t}$ in Eqn. (3) with predicted velocity $v_{\theta,t}$ using the relation $\epsilon_{\theta,t}=x_{t}+\alpha_{t}v_{\theta,t}$. Additionally, we substitute $\sqrt{\bar{\alpha}_{t}}$ and $\sqrt{1-\bar{\alpha}_{t}}$ in Eqn. (3) and Eqn. (4) with $\alpha^{\prime}_{t}$ and $\sigma^{\prime}_{t}$ in the forward process of Flow Matching, i.e., $x_{t}=\alpha^{\prime}_{t}x_{0}+\sigma^{\prime}_{t}\epsilon$. Then the backward process can be reformulated as:
$x_{t-1}=\alpha^{\prime}_{t-1}(x_{t}-\sigma^{\prime}_{t}v_{\theta,t})+\sqrt{(\sigma^{\prime}_{t-1})^{2}-\sigma_{t}^{2}}(\alpha^{\prime}_{t}v_{\theta,t}+x_{t})+\sigma_{t}\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (14)
$\sigma_{t}=\eta\sqrt{\left(\frac{\sigma^{\prime}_{t-1}}{\sigma^{\prime}_{t}}\right)^{2}\left(1-\left(\frac{\alpha^{\prime}_{t}}{\alpha^{\prime}_{t-1}}\right)^{2}\right)}.$ (15)
Finally, we employ LPO on SD3-medium. As shown in Tab. 14 and Tab. 14, even though the training data Pick-a-Pic only contains images from SDXL and SD2.1, LPO still demonstrates notable performance gains on various metrics. We believe that if there is higher-quality training data, such as that in InterComp, the performance gain will be more pronounced.
Experimental Results on HPDv2. To validate the generalization of LRM and LPO, we conduct experiments on the HPDv2 [17] dataset without any hyperparameter optimization. Specifically, we train LRM on the entire HPDv2 training set and then randomly sample 4,000 prompts from this set for LPO training. In addition to the previous evaluation metrics, we also randomly sample 400 prompts from the HPDv2 test set and calculate the general and aesthetic scores for these prompts. As shown in Tab. 17, Tab. 17, and Tab. 17, LPO significantly outperforms the original models across various evaluation dimensions.
Table 15: Performance of LPO on HPDv2. General and aesthetic preference scores on Pick-a-Pic validation unique set, along with GenEval [40] scores. If not specified, we use 20 inference steps.
Model Method Dataset PickScore ImageReward HPSv2 HPSv2.1 Aesthetic GenEval (20 Step) GenEval (50 Step)
SD1.5 [14] Original - 20.56 0.0076 26.46 24.05 5.468 42.56 42.53
LPO HPDv2 [17] 21.24 0.7248 28.13 28.93 5.917 47.29 48.27
Table 16: Performance of LPO on HPDv2. General and aesthetic preference scores on 400 randomly sampled HPDv2 test prompts with 20 inference steps.
Model Method Dataset PickScore ImageReward HPSv2 HPSv2.1 Aesthetic
SD1.5 [14] Original - 20.56 0.0076 26.46 24.05 5.468
LPO HPDv2 [17] 21.61 0.7282 28.81 29.25 6.019
Table 17: Performance of LPO on HPDv2. Quantitative results on T2I-CompBench++ [41] with 20 inference steps. LPO is trained on the HPDv2 [17] dataset.
SD1.5 [14] Original 0.3783 0.3616 0.4172 0.1230 0.2967 0.4485 0.3104 0.2999
LPO 0.5761 0.5067 0.6013 0.1671 0.3345 0.4671 0.3035 0.3332
@@
Specific Reward Model (VLM-Based) Aesthetic [38] - - - 54.03
CLIP Score (CLIP-H) [9] - - - 61.84
VQAScore [39] - - - 59.16
General Reward Model (VLM-Based) ImageReward [18] 0.108 0.425 0.417 62.66
HPSv2 [17] 0.007 0.602 0.406 64.76
HPSv2.1 [17] 0.191 0.432 0.332 65.58
PickScore [16] 0.066 0.490 0.402 71.93
@@
LRM as Pixel-Wise Reward Model. The LRM can also serve as a pixel-wise reward model when combined with the corresponding VAE encoder. Tab. 18 shows the correlations of different reward models with aesthetics and text-image alignment. Aes-Corr is employed to assess the correlation with aesthetics, while the CLIP-Corr and VQA-Corr are utilized to measure alignment correlation. The calculation method for VQA-Corr is similar to that of CLIP-Corr, as described in Sec. 5.1, with the CLIP Score replaced by VQAScore. It is observed that the LRMs, especially LRM-XL, exhibit aesthetic and alignment correlations comparable to those of VLM-based models, as indicated by similar Aes-Corr and VQA-Corr values. The CLIP-Corr of LRMs is slightly lower than the average value of VLM-based reward models, likely because VLM-based models are fine-tuned versions of CLIP [9] or BLIP [19], leading to higher similarity to CLIP. The accuracy on validation and test sets of Pick-a-Pic is also present in Tab. 18. LRMs exhibit competitive performance, with larger variants (LRM-XL) achieving higher accuracy. Moreover, we observe that adding slight noise to the latent images ($t=200$) leads to improved accuracy, which may be attributed to the fact that slight noise can better activate the feature extraction capabilities and improve the generalization of diffusion models.
Appendix C Additional Ablation Experiments
Table 19: Comparison between SPO and LPO without MPCF. The general and aesthetic scores are calculated on the Pick-a-Pic validation unique set. If not specified, we use 20 inference steps.
Model Method PickScore ImageReward HPSv2 HPSv2.1 Aesthetic GenEval (20 Step) GenEval (50 Step)
SD1.5 [14] SPO 21.22 0.1678 26.73 25.83 5.927 40.46 41.53
LPO wo MPCF 21.49 0.4406 27.39 27.61 5.772 45.66 46.21
Table 20: Comparison results on T2I-CompBench++ [41] with 20 inference steps between SPO and LPO without MPCF.
SD1.5 [14] SPO 0.4112 0.4019 0.4044 0.1301 0.2909 0.4372 0.3008 0.2988
LPO 0.4717 0.4002 0.4846 0.1793 0.3460 0.4596 0.3094 0.3171
Comparison between SPO and LPO without MPCF. We conduct a comparison between SPO and LPO on identical training data. Specifically, we compare SPO with LPO without the MPCF strategy. As illustrated in Tab. 20 and Tab. 20, LPO without MPCF still outperforms SPO, especially in general and text-image alignment preference, demonstrating the effectiveness of reward modeling and preference optimization in the noisy latent space.
Table 21: Comparison between SPO and LPO with different training timestep ranges. We use 20 inference steps. denotes the results quoted from [8].
Model Method Range of $t$ PickScore ImageReward HPSv2 HPSv2.1 Aesthetic GenEval
SD1.5 [14] original - 20.56 0.0076 26.46 24.05 5.468 42.56
SPO [0, 750] 21.43 0.1712 26.45 - 5.887 -
[0, 950] 19.77 -0.4529 22.72 - 5.111
LPO [0, 700] 21.02 0.3087 27.10 26.25 5.765 44.93
[750, 950] 21.54 0.6337 27.47 27.64 5.853 48.28
[0, 950] 21.69 0.6588 27.64 27.86 5.945 48.39
Table 22: Comparison of timestep inputs of LRM during LPO sampling. We use 20 inference steps.
Model Input Timestep PickScore ImageReward HPSv2 HPSv2.1 Aesthetic GenEval
SD1.5 [14] Real Timestep 21.69 0.6588 27.64 27.86 5.945 48.39
0 19.64 -0.5043 25.37 21.74 5.283 32.21